OpenStudy (shayaan_mustafa):

Find slope of the tangent line to the parametric curve \[x=t^2+1\space\space,\space\space y=\frac{t}{2}\space\text{ at} \space \space t=-1 \text{ and at} \space \space t=1 \text{ by eliminating parameters}\]

OpenStudy (shayaan_mustafa):

I know how to do it by without eliminating parameters. I want to know the method of elimination of parameters.

OpenStudy (turingtest):

just solve both equations for t *sorry my connection is bad right now

OpenStudy (shayaan_mustafa):

Why both? Solve one and substitute that value of t in another equation. huh?

OpenStudy (turingtest):

\[t=\sqrt{x-1}=2y\implies y=\frac12\sqrt{x-1}\]at\[t=-1\implies x=2\]so we now want\[\frac{dy}{dx}|_{x=2}\]

OpenStudy (shayaan_mustafa):

Sorry it is not square.

OpenStudy (shayaan_mustafa):

\[\large y^2=\pm\frac{\sqrt{x-1}}{2}\] This is what I got.

OpenStudy (turingtest):

we can only do one half of the parabola at a time, so you have two equations for the tangent lines

OpenStudy (shayaan_mustafa):

OK at t=1 and at t=-1 I have x=2 for both. And I have equation look like this\[\large y=\pm\frac{\sqrt{x-1}}{2}\]Which one can I use equation with + sign or equation with -sign? And also in this question we have t=1 and t=-1 both are giving same value x=2. What to do when I have two different values of x's on different t's, which one I have to choose then?

OpenStudy (amistre64):

r(t) = <x(t),y(t)> r'(t) = <x'(t),y'(t)> right?

OpenStudy (shayaan_mustafa):

yes.

OpenStudy (amistre64):

then r'(t) = <2t, 1/2>; at t=1 we have a tangent vector of <2,1/2> then

OpenStudy (amistre64):

so what is it we are trying to accomplish here?

OpenStudy (shayaan_mustafa):

slope of tangent line.

OpenStudy (shayaan_mustafa):

Wait wait.. I think I got now.

OpenStudy (amistre64):

slope = y/x 1/2 --- = 1/4 2

OpenStudy (shayaan_mustafa):

Yes right.

OpenStudy (shayaan_mustafa):

At t=-1?

OpenStudy (shayaan_mustafa):

-1/4 right?

OpenStudy (amistre64):

1/2 ---- is our general set up for this; so when t=-1, -1/4 2t

OpenStudy (shayaan_mustafa):

Yeah.. Thanks.. I want to ask you one thing. You used vector here?

OpenStudy (shayaan_mustafa):

In book nothing given as <> (with these brackets) . But you use these brackets. So that's why I am asking to you.

OpenStudy (amistre64):

i did use vector yes; since a parametric defines a vector in terms of x and y components

OpenStudy (shayaan_mustafa):

OK..

OpenStudy (amistre64):

if i were to define it another way: x = t^2 + 1 ; y = t/2 t = 2y x = (2y)^2 + 1 x = 4y^2 + 1 ; derive it implicitly 1 = 8y y' y' = 1/8y ; y = t/2 y' = 1/(8t/2) = 1/4t

OpenStudy (shayaan_mustafa):

Yeah Thanks.. Now it is cleared.

OpenStudy (shayaan_mustafa):

@TuringTest I thank you too friend. :)

OpenStudy (turingtest):

welcome, but amistre's answer was better :)